# augmented reality AP average precision

### v1

@@
Hannah Schieber1, Dominik Frischmann1, Simon Boche3, Victor Schaack1, Angela Schoellig2,
@@
1Hannah Schieber, Dominik Frischmann, Victor Schaack, and Daniel Roth are with the Technical University of Munich, Human-Centered Computing and Extended Reality Lab, TUM University Hospital, Clinic for Orthopedics and Sports Orthopedics, Munich Institute of Robotics and Machine Intelligence (MIRMI) hannah.schieber@tum.de, daniel.roth@tum.de2Angela Schoellig is with Technical University of Munich angela.schoellig@tum.de2Simon Boche is with Technical University of Munich simon.boche@tum.de2Stefan Leutenegger is with the Mobile Robotics Lab, Department of Mechanical and Process Engineering, ETH Zurich lestefan@ethz.ch
@@ Abstract
Mobile reconstruction for autonomous aerial robotics holds strong potential for critical applications such as tele-guidance and disaster response. These tasks demand both accurate 3D reconstruction and fast scene processing. Instead of reconstructing the entire scene in detail, it is often more efficient to focus on specific objects, i.e., points of interest (PoIs). Mobile robots equipped with advanced sensing can usually detect these early during data acquisition or preliminary analysis, reducing the need for full-scene optimization. Gaussian Splatting (GS) has recently shown promise in delivering high-quality novel view synthesis and 3D representation by an incremental learning process. Extending GS with scene editing, semantics adds useful per-splat features to isolate objects effectively.
Semantic 3D Gaussian editing can already be achieved before the full training cycle is completed, reducing the overall training time. Moreover, the semantically relevant area, the PoI, is usually already known during capturing. To balance high-quality reconstruction with reduced training time, we propose CoRe-GS. We first generate a coarse segmentation-ready scene with semantic GS and then refine it for the semantic object using our novel color-based effective filtering for effective object isolation. This is speeding up the training process to be about a quarter less than a full training cycle for semantic GS. We evaluate our approach on two datasets, SCRREAM (real-world, outdoor) and NeRDS 360 (synthetic, indoor), showing reduced runtime and higher novel-view-synthesis quality.
@@ I INTRODUCTION
Capturing and reconstructing our world is beneficial for a wide range of domains [1, 2, 3, 4], particularly when enhanced beyond traditional 2D imaging. By enabling 3D reconstructions, such captures offer more detailed and explorable scene representations. Applications span aerial imaging [5], medicine [6, 7], cultural heritage preservation and tourism [2, 3, 4], architecture [8], construction [9], planning [10], and robotic manipulation [11, 12, 13]. High-fidelity 3D scene representations are essential for simulation environments, remote mission planning, and robotic navigation, especially in drone- and robot-captured settings. In the context of novel view synthesis (NVS), Gaussian Splatting (GS) [14] presents an efficient approach with a trade-off between training time and scene quality.
Figure 1: Comparison of traditional semantic GS following the definition of Ye et al. [15] compared to our approach leveraging directly the knowledge of specific PoIs during training time.
@@
To further optimize such pipelines, especially in drone-based visual captures, PoI selection can be a valuable strategy to focus on relevant scene areas. A PoI typically guides the camera toward a specific object or region during capture, ensuring that this area is well-represented in the resulting dataset. When constructing 3D representations from these images, PoI metadata can continue to play a role, informing downstream semantic tasks, guiding training prioritization, or supporting scene understanding. While learning-based 3D representations often require significant training overhead, integrating PoI awareness can streamline this process. In particular, semantics tied to PoIs can drive efficient scene representation, offering targeted improvements. Currently, semantic approaches allowing the extraction of PoI objects often suffer from distracting scene parts (floaters) in the filtered scene representation.
@@
For scene refinement, active view selection [16, 17] is a standard methodology. However, these approaches typically optimize the entire scene. Optimizing specific scene parts is usually not applied, and post-processing approaches suffer from outliers (floaters) [4, 18, 15]. Optimizing specific scene parts can be acquired with semantics, but incorporating semantic segmentation often introduces additional computational overhead. Although semantic GS offers sufficient fidelity to support selective refinement of drone-captured regions, it is generally applied to the whole scene [4, 15]. This extends the overall runtime for initial reconstruction.
@@
To address this challenge, we propose a refinement of PoIs in GS. Our method reduces training time for high-quality scene reconstruction by focusing on a single PoI. We extend classical and semantic GS by integrating PoI extraction, accelerating the overall training process. Once a segmentation-ready scene (coarse scene) is created, further processing refines only the extracted object, the PoI, in roughly a quarter of the time required by standard semantic GS [15]. By effectively mitigating color space discrepancies during PoI refinement, our approach enhances the visual quality of the resulting scenes. We demonstrate our method in both indoor and outdoor settings using scenes from the SCRREAM [19] and NeRDS 360 [20] datasets.
@@
  • A PoI refinement for an efficient and more accurate editable GS.
  • A color-based effective filtering method, for GS PoI refinement.
@@ II RELATED WORK
Our approach targets editable radiance fields and reduces the overall training time to extract PoIs from these radiance field representations. Therefore, we provide an overview of state-of-the-art editing of radiance fields, refinement, and semantics.
@@
While enhancing the visual quality of the scene representation is one approach, Peralta et al. [21] focus on adapting the capturing path. Similarly, Chen et al. [22] introduce GenNBV, which optimizes drone camera paths using reinforcement learning. While these approaches apply to a broader scope, others directly target GS optimization. Aside from optimizing the capturing path, in already captured scenes, FisherRF [17] applies uncertainty quantification to select ideal views within GS. Lit et al. [16] leverages efficient viewpoint sampling and path planning combined with a dense scene representation via GS.
Refinement of the scene based on semantic PoI offers a different solution to the problem, as the complete scene does not need to be refined.
II-B Semantic and Language-Guided Radiance Fields
Editable radiance fields often employ semantics or language guidance to adapt the final scene representation [23, 15, 4, 24, 25, 26]. Different final applications can be considered, ranging from object removal to search until object extraction.
@@ II-B1 Object Removal
Although GS [14], now dominates radiance fields, its predecessor, neural radiance fields (NeRF[27, 28], already presented editable approaches [29, 28]. Inpainting can be used to remove objects from the initial scene or from the learned scene representation [30, 31, 32] with the goal to keep an accurate geometry of the scene. Yin et al. [30] leveraged SAM [33] and with LaMa [34] to inpaint selected objects within a NeRF. In a similar approach, Weder et al. [31] proposed an inpainting-based NeRF solution designed to remove unwanted objects. Similarly, GS has been addressing object removal tasks using segmentation masks first of the unwanted object [32]. For GS, Huang et al. address this with SAM masks and depth-guided inpainting.
Many GS approaches allow object removal [23, 15] and, combined with an additional inpainting model like LaMa [34], also enable inpainted Gaussians.
@@ II-B2 Semantic Search and Editing
While NeRF already showed great potential, GS [14] offers better rendering capability and adaptability, due to its explicit scene representation. Its real-time rendering capability is especially beneficial for scene exploration. Often, multiple models are combined. Yu et al. [25] leverage CLIP, Droid-SLAM [35], and GS for robotic navigation. In addition to language guidance, HAMMER [36] produces semantic GS maps from multiple input devices, such as robots and smart glasses, allowing direct editing of the resulting scene representations.
Similarly, Qu et al. [37] proposed a point-based editing approach that enables manipulation of Gaussians directly within the scene. Furthermore, segmentation-based GS allows scene editing per class, often proposed as a downstream task [23, 15]. Ye et al. [15] proposed segmentation-based grouping, with the so-called Gaussian Grouping (GG), extending the Gaussian representation with a feature field for classifier-based segmentation. Moreover, they demonstrated scene editing, such as inpainting and object removal, as a downstream task. Within GAGA [23], Lyu et al. [23] proposed segmentation-based GS using a memory bank. Although segmentation quality can be improved with GAGA, training time also increases, as the approach uses pure GS [14] as input, followed by several semantic training and label preparation steps.
While many approaches [15, 23] leverage foundation models [38, 33] for segmentation ground truth, Schieber et al. [4] use a supervised approach, effectively removing large object parts in outdoor scenes. While it is promising for large objects, foundation models have broader scalability to unknown classes.
@@ II-C Applied Editable Radiance Fields
Adding semantic segmentation to the radiance field representation provides the opportunity to later edit the explicit GS scene representation. Many works [23, 15, 4], demonstrate this ability in a post-processing step, others directly integrate it into applications. Li et al. [11] added semantics for robotic manipulation for a limited number of views. They address language-guided grasping. Another grasping solution is proposed by Ji et al. [12]. They leverage GS, MobileSAM, and CLIP to propose potential object grasp positions.
@@ II-D Research Gap
While scene editing [15, 4, 26], path planning [21, 22], and camera view selection [17] are widely explored, optimizing the semantic representations used within these contexts can still be optimized. Complete semantic training is costly, and post-processing PoI extraction suffers from inconsistency (floaters). To address these challenges, we propose PoI refinement GS, where a coarse semantic GS provides initial scene understanding and PoI are selectively refined, achieving high quality with reduced training time.
@@ III-A Gaussian Segmentation and Point-of-Interest Extraction
We apply the training process of semantic GS [15] for 3k iterations and obtain a coarse scene representation, see Figure 2. The pre-segmentation step used enables each part of the scene to be represented by a semantic class or instance class, depending on the given database. Using the assigned class ID, a filtering step is performed to extract the PoI.
Figure 2: Our approach. Our approach first trains a semantic GS scene representation (left) [15] for a set of initial iterations (3k). Followed by that, the PoI is extracted based on the given segmentation class from the semantic GS (center). Based on the segmentation class ID, we extract only the Gaussians per PoI. In a last step, the PoI is refined using our periodic scene filtering with the previously extracted furthest color.
During the pre-extraction phase, the PoI is selected based on its class ID, such as the class “car”, see Figure 2. The images are filtered on the basis of the presence of this class. The Gaussians are then filtered accordingly, preserving only those that predominantly belong to the selected PoI.
For each retained image, binary masks ($M$) are generated to isolate the target object: for each pixel ($u,v$) corresponding to the segmentation class are assigned a value of $1$ is assigned, while all other pixels are set to 0:
$M(u,v)=\begin{cases}1,&\text{if }(u,v)\text{ belongs to the target object ID}\\ 0,&\text{otherwise}\end{cases}$ (1)
These binary masks ($M$) are used to effectively remove additional information. This is required in the later refinement process to ensure that only the relevant parts of the scene within the PoI are processed in the refinement step.
In addition, the color distribution of the image is analyzed for each selected image to determine the most pronounced color in the image in relation to the other colors in the image. To speed up this process, the images are downscaled by a factor of $0.5$. Unique colors are then extracted from the downscaled images, and a KD-tree is created based on these colors. The set of existing image colors is given by:
@@
is sampled systematically across the RGB spectrum ($r,g,b$). $P$ is queried against the constructed KD-tree. The color with the greatest minimum Euclidean distance ($\|p_{j}-c_{i}\|_{2}$) from all existing image colors is identified as the “furthest” color:
@@
where $r,g$, and $b$ represent the RGB values normalized to the interval $[0,1]$. This furthest color ($p^{*}$), along with its corresponding maximum minimum distance value, is recorded for each image. These results are subsequently saved for analysis and utilized during the refinement process.
@@ III-B Point of Interest Refinement
Based on extracted colors, the frequency of the most common color is analyzed. The most frequent color is used as the universal background for the subsequent rendering. To focus refinement on the selected PoI, the training images and loss calculations are masked with binary masks $M(x,y)$.
We initialize our GS using the filtered points of the PoI and start the training process for the remaining 27k iterations. Additionally, our GS is periodically filtered to eliminate potential artifacts during refinement.
@@ III-B1 Periodic Scene Filtering
While segmentation classes can initially be utilized for isolating the PoI due to the semantic information provided by any 3D semantic GS, such segmentation or instance classes are not leveraged during the refinement step, as the semantic rasterizer and classifier lead to an increase in training time. Instead, we employ our color-based filtering strategy for additional pruning and adjustment of the PoI GS during refinement. In the pre-extraction step, we received the most distinct (“furthest”) color ($p^{*}$) relative to the colors already present, measured by Euclidean distance in RGB color space.
Using $p^{*}$, we effectively filter out unwanted Gaussians. Given the substantial color difference between $p^{*}$ and meaningful Gaussians, unwanted artifacts can be identified and removed through color-based filtering. Specifically, for each view, the color of each Gaussian is computed by evaluating its spherical harmonics (SH) coefficients relative to the camera’s viewing direction. The resulting RGB color is then compared to the previously identified $p^{*}$. The Euclidean distance between the Gaussians computed RGB color and $p^{*}$ is calculated. If this distance falls below a threshold, the Gaussian is considered an artifact and marked for removal. All Gaussians identified as artifacts based on this criterion are pruned from the scene representation, thereby significantly reducing unwanted visual artifacts that typically manifest around object boundaries. Given the average distance $d_{avg}$ computed earlier from the $p^{*}$ selection process, the Gaussians selected for removal at a removal iteration are those satisfying:
$G_{\text{remove}}=\{G_{i}\mid d(c_{G_{i}},c_{\text{bg}})<d_{\text{remove}}\}$ (6)
@@
The removal threshold defaults to $t_{r}=0.5$. This process ensures that the refinement specifically targets meaningful regions and maintains clean, artifact-free reconstructions throughout iterative refinement.

### v2

@@
Hannah Schieber1, Dominik Frischmann1, Victor Schaack1, Simon Boche3, Angela Schoellig3,
@@
1Hannah Schieber, Dominik Frischmann, Victor Schaack, and Daniel Roth are with the Technical University of Munich, Human-Centered Computing and Extended Reality Lab, TUM University Hospital, Clinic for Orthopedics and Sports Orthopedics, Munich Institute of Robotics and Machine Intelligence (MIRMI) hannah.schieber@tum.de, daniel.roth@tum.de2Simon Boche is with Technical University of Munich simon.boche@tum.de3Angela Schoellig is with Technical University of Munich angela.schoellig@tum.de2Stefan Leutenegger is with the Mobile Robotics Lab, Department of Mechanical and Process Engineering, ETH Zurich lestefan@ethz.ch
@@ Abstract
Mobile reconstruction has the potential to support time-critical tasks such as tele-guidance and disaster response, where operators must quickly gain an accurate understanding of the environment. Full high-fidelity scene reconstruction is computationally expensive and often unnecessary when only specific points of interest (POIs) matter for timely decision making. We address this challenge with CoRe-GS, a semantic POI-focused extension of Gaussian Splatting (GS). Instead of optimizing every scene element uniformly, CoRe-GS first produces a fast segmentation-ready GS representation and then selectively refines splats belonging to semantically relevant POIs detected during data acquisition. This targeted refinement reduces training time to 25% compared to full semantic GS while improving novel view synthesis quality in the areas that matter most. We validate CoRe-GS on both real-world (SCRREAM) and synthetic (NeRDS 360) datasets, demonstrating that prioritizing POIs enables faster and higher-quality mobile reconstruction tailored to operational needs.
@@ I INTRODUCTION
Capturing and reconstructing real-world environments in 3D is increasingly essential for domains ranging from aerial imaging and disaster response to medicine, cultural heritage, and robotic navigation [1, 2, 3, 4, 5, 6]. 3D representations, like Gaussian Splatting (GS) [7] provide detailed, explorable models that support simulation, remote mission planning and operator decision making, especially in drone and robot captures. In disaster response, rapid situational awareness is critical because teams often need actionable 3D maps within minutes rather than hours. Full-scene reconstruction can be computationally costly, time-consuming, and unnecessary when only specific POI are operationally relevant. Focusing reconstruction on mission-critical regions rather than entire environments cuts processing time and memory requirements and enables faster deployment and decision-making in time-sensitive operations.
@@
To further optimize such pipelines, especially in drone-based visual captures, Point of Interest (POI) selection can be a valuable strategy to focus on relevant areas. During capture, a POI typically guides the camera towards a specific object or region. When constructing 3D representations from these images, POI metadata can continue to play a role, informing downstream semantic segmentation tasks, guiding training prioritization, or supporting scene understanding. While learning-based 3D representations of the entire scene often require significant training overhead, integrating POI awareness can streamline this process. Currently, semantic approaches allowing the extraction of POI objects often suffer from outliers (floaters), see Fig. 1.
Figure 1: Comparison of traditional semantic GS following Ye et al. [8] compared to our CoRe-GS leveraging directly the knowledge of specific POIs during training time.
@@
For scene refinement, active view selection [9, 10] is the standard method. However, these approaches typically optimize the entire scene. Optimizing only specific parts is usually not done, and post-processing approaches suffer from outliers (floaters) [4, 11, 8]. Semantic segmentation can be used to optimize particular parts, but incorporating these often introduces computational overhead. Although semantic GS offers sufficient fidelity to support selective refinement, it is generally applied to the entire scene [4, 8, 12]. This extends the overall runtime for initial reconstruction.
@@
To address this challenge, we propose CoRe-GS, a refinement of POIs in GS. CoRe-GS reduces training time for high-quality scene reconstruction by focusing on a single POI. We extend classical and semantic GS by integrating POI extraction, accelerating the overall training process. Once a segmentation-ready scene is created, further processing refines only the extracted object, the POI, in roughly a quarter of the time required by standard semantic GS [8]. By effectively mitigating color space discrepancies during POI refinement, our approach enhances the visual quality of the resulting scenes. We demonstrate our method in both indoor and outdoor settings using scenes from the SCRREAM [13] and NeRDS 360 [14] datasets.
@@
  • A novel POI refinement for semantic object focus and efficient editable GS.
  • A novel color-based effective filtering method, for GS POI refinement, addressing outliers.
@@ II RELATED WORK
While enhancing the visual quality of the scene representation is one approach, Peralta et al. [15] focus on adapting the capturing path. Similarly, Chen et al. [16] introduce GenNBV, which optimizes drone camera paths using reinforcement learning. While these approaches apply to a broader scope, others directly target GS optimization [10]. FisherRF [10] applies uncertainty quantification to select ideal views within GS. Li et al. [9] leverage efficient viewpoint sampling and path planning combined with a dense scene representation via GS.
Refinement of the scene based on semantic POIs improves efficiency in another direction than optimizing the initial capturing path. With POIs-based refinement, only the relevant parts are updated leading to a runtime improvement.
II-B Editable Radiance Fields
Editable radiance fields often employ semantics or language guidance to adapt the final scene representation [12, 8, 4, 17, 18, 19]. Different applications can be considered, ranging from object removal /extraction to object search or inpainting.
@@ II-B1 Object Removal
Although GS [7] now dominates radiance fields, its predecessor neural radiance fields (NeRF[20, 21, 22] already offered editable approaches [21, 22]. Inpainting can remove objects from the learned radiance field using an inpainting training step [23, 24, 25]. Yin et al. [23] leveraged SAM [26] and LaMa [27] to inpaint selected objectsfor NeRF. In a similar approach, Weder et al. [24] proposed an inpainting-based NeRF solution to remove unwanted objects. Similarly, object removal tasks have been addressed in GS using segmentation masks, for example, Huang et al. [25], address this with SAM and depth-guided inpainting. In addition, semantic GS approaches provide another way to handle object removal [12, 8] and, when combined with an inpainting model such as LaMa [27], enable inpainting of Gaussians.
@@ II-B2 Semantic Search and Editing
Although NeRF already showed great potential, GS [7] offers better rendering capability and adaptability due to its explicit scene representation. Often, multiple models are combined. Yu et al. [18] leverage CLIP, Droid-SLAM [28], and GS for robotic navigation. In addition to language guidance, HAMMER [29] produces semantic GS maps from multiple input devices, such as robots and smart glasses, allowing for direct editing of the resulting scene representations.
Similarly, Qu et al. [30] proposed a point-based editing approach that enables the manipulation of Gaussians directly within the scene. Furthermore, segmentation-based GS allows scene editing per class, often proposed as a downstream task [8, 12]. Ye et al. [8] proposed segmentation-based grouping, with the so-called Gaussian Grouping (GG), extending the Gaussian representation with a feature field for classifier-based segmentation. Moreover, they demonstrated scene editing, such as inpainting and object removal, as a downstream task. Within GAGA, Lyu et al. [12] proposed segmentation-based GS using a memory bank. Although segmentation quality can be improved with GAGA, training time also increases, as the approach uses pure GS [7] as input, followed by several semantic training and label preparation steps. For large objects and editing large scenes, Schieber et al. [4] show that a supervised model can remove objects more effectively compared to foundation models. Although this is promising for large objects, foundation models offer broader scalability for unknown classes.
@@ II-C Applied Editable Radiance Fields
Adding semantic segmentation to the radiance field representation provides the opportunity to later edit the explicit GS scene representation. Semantic GS approaches [12, 8, 4] demonstrate this ability in a post-processing step. Others directly integrate it into robotic applications [31, 32]. Li et al. [32] added semantics for robotic manipulation for a limited number of views. They address language-guided grasping. Another grasping solution is proposed by Ji et al. [31]. They leverage GS, MobileSAM, and CLIP to propose potential object grasp positions.
@@ II-D Research Gap
Scene editing [8, 4, 19], path planning [15, 16], and camera selection [10] have already been extensively researched, but there is still room for improvement in optimizing the semantic representations used in these contexts. Complete semantic training is costly, and subsequent POI extraction often leads to inconsistencies (e.g., floaters). We introduce CoRe-GS, a POI refinement approach GS that uses a segmentation-ready GS for initial understanding and selectively refines POI, achieving high-quality results with significantly reduced training time.
Figure 2: Our approach first trains GG [8] (initial segmentation-ready GS) for 3k initial iterations. Followed by that, the POI is extracted based on the given segmentation class (POI selection). In a last step, the POI is refined using our periodic scene filtering with the previously extracted furthest color (mask and color extraction and POI refinement).
@@ III-A Gaussian Segmentation and Point-of-Interest Extraction
We apply the training process of semantic GS [8] for 3k iterations and obtain a coarse, segmentation-ready scene representation, see Figure 2. The pre-segmentation step enables each part of the scene to be represented by a semantic class or instance class, depending on the given database. During pre-extraction, POIs are selected by class ID (e.g., “car”, see Fig. 2). Images containing the chosen class are retained, and only Gaussians predominantly associated with the selected POI are preserved.
For each retained image, binary masks ($M$) are generated to isolate the target object: for each pixel ($u,v$) corresponding to the segmentation class, a value of $1$ is assigned, while all other pixels are set to 0. These binary masks ($M$) are used to effectively remove additional information. This ensures, in the later refinement process, that only the relevant parts of the scene within the POI are processed in the refinement step. During training, GS often produces floaters, unwanted Gaussians, in the background. Relying solely on a masked loss during refinement leads to soft or fuzzy contours, as areas outside the POI are effectively ignored. Alternatively, replacing the background with a generic color like black or white can result in floaters adopting those common colors, which are also likely to appear in the foreground object. This makes it difficult to distinguish and filter them later. To address both issues, we introduce a color filtering step that assigns a background color maximally distinct from the scene content, enabling effective removal of floaters based on color distance.
To employ it, we analyze the color distribution of the image. For each image, we determine the most distant color from the colors present in the image. We speed up this process by downscaling the images by a factor of $0.5$. Unique furthest colors are then extracted from the downscaled images, and a KD-tree is created based on these furthest colors. The set of existing image colors is given by:
@@
is sampled systematically across the RGB spectrum ($r,g,b$). $P$ is queried against the constructed KD-tree. The color with the greatest minimum Euclidean distance ($\|p_{j}-c_{i}\|_{2}$) from all existing image colors is identified as the “furthest” color $p^{*}$, and later used for rendering, to color floaters uniquely:
@@
where $r,g$, and $b$ represent the RGB values normalized to the interval $[0,1]$. This information is utilized during the refinement process to render the background color and to identify unwanted background floaters.
Figure 3: Exemplary occlusion removal comparison on the scene SF_6thAndMission_medium6 (bottom) and SF_6thAndMission_medium10 (top) after 30,000 iterations. Post-training isolation using $G$ partially recovers occlusions, with “convex hull” usage influencing the result. Our approach produces a clean background.
@@ III-B Point of Interest Refinement
The “furthest” color is used as a universal background for the subsequent rendering. To focus refinement on the selected POI, the training images and loss calculations are masked with binary masks $M(x,y)$. We initialize our POI refinement in CoRe-GS using the filtered points of the POI and start the training process for the remaining 27,000 iterations. In addition, our GS is regularly filtered using the “furthest” color to eliminate potential floaters during refinement.
@@ III-B1 Periodic Scene Filtering
Although segmentation classes from a 3D semantic GS can initially isolate the POI, they are not used during refinement because the semantic rasterizer and classifier significantly increase training time. Instead, we employ our color-based filtering strategy for additional pruning and adjustment of the POI GS. In the pre-extraction step, we received the most distinct (“furthest”) color ($p^{*}$) relative to the colors already present, measured by Euclidean distance in RGB color space ($d_{avg}$).
Using $p^{*}$, we effectively filter out unwanted Gaussians. Given the substantial color difference between $p^{*}$ and meaningful Gaussians, unwanted artifacts (floaters) can be identified and removed through color-based filtering. Specifically, for each view, the color of each Gaussian is computed by evaluating its spherical harmonics (SH) coefficients relative to the camera’s viewing direction. The resulting RGB color is then compared to the previously identified ($p^{*}$). Then the Euclidean distance between the Gaussians computed RGB color and $p^{*}$ is calculated. If this distance falls below a threshold, the Gaussian is considered an artifact and marked for removal. All Gaussians identified as artifacts based on this criterion are pruned from the scene representation, thereby significantly reducing unwanted visual artifacts that typically manifest around object boundaries. Given the average distance ($d_{avg}$) computed earlier from selecting $P$, the Gaussians ($G$) selected for removal $G_{\text{remove}}$ at a removal iteration are those satisfying:
$G_{\text{remove}}=\{G_{i}\mid d(c_{G_{i}},p^{*})<d_{\text{remove}}\}$ (5)
@@
The removal threshold defaults to $t_{r}=0.5$, a value we determined empirically. This process ensures that the refinement specifically targets meaningful splats and maintains clean, artifact-free refinement.